OpenStudy (anonymous):

How many kilojoules of energy are absorbed when 800.0g of water is heated from 20.0º C to 50.0º C?

OpenStudy (anonymous):

q=cm(T1-T2) q=heat(energy) c=heat capacity (equals 1 for water) T=Temp change q=1cal/g°C*800g*30°C q=24000cal But in kcal, it would be 24kcal